2x+4y=-10;(4,2)
write equation of a line that is perpendicular to the given line and that passes through the given point

Answer by robertb(5830)

The unknown line would have the form
4x-2y = c, where c is an undetermined constant. To find c use the coordinates of the given point:
4*4 - 2*2 = 16 - 4 = 12.
Therefore the equation of the line is
4x-2y = 12.